# TIPO: Text to Image with Text Presampling for Prompt Optimization

### v2

@@ Abstract
TIPO (Text to Image with text pre-sampling for Prompt Optimization) is an innovative framework designed to enhance text-to-image (T2I) generation by language model (LM) for automatic prompt engineering. By refining and extending user-provided prompts, TIPO bridges the gap between simple inputs and the detailed prompts required for high-quality image generation. Unlike previous approaches that rely on Large Language Models (LLMs) or reinforcement learning (RL), TIPO adjusts user input prompts with the distribution of a trained prompt dataset, eliminating the need for complex runtime cost via lightweight model. This pre-sampling approach enables efficient and scalable prompt optimization. Experimental results demonstrate TIPO’s effectiveness in improving aesthetic scores, reducing image corruption, and better aligning generated images with dataset distributions. These findings highlight the critical role of prompt engineering in T2I systems and open avenues for broader applications of automatic prompt refinement.
@@ 1 Introduction
Recent advancements in Text-to-Image (T2I) generative models have revolutionized creative applications [47, 7, 21, 55, 51, 52, 57, 53, 48, 56, 15, 14, 34, 22, 8]. These models tend to perform better when provided with longer, more detailed prompts, which describes specific elements like style, composition, and context. However, this reliance on highly descriptive inputs is a significant bottleneck for generating high-quality images, especially for users who prefer or require simpler inputs. The need for intricate, nuanced prompts often results in a higher barrier to entry, limiting accessibility and ease of use for those unfamiliar with the complexities of prompt engineering.
Previous works have attempted to extend prompts with optimization through various methods. Some approaches use LLMs to expand prompts [41]; however, due to the misalignment between the LLM’s text-rich training corpus and the T2I model’s caption-focused corpus, only a limited set of manually-engineered prompts yield effective results. Other methods rely on curated prompt sets [5, 17], which, while potentially incorporating user preferences, fail to fully leverage the T2I model’s trained prompt distribution. Additionally, reinforcement learning-based techniques [26] require multi-turn inference, incurring high computational costs, and often struggle to generate diverse outputs due to the model’s inherent corpus limitations.
To address these challenges, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. TIPO introduces ‘text pre-sampling’ preprocess procedure in inference pipeline, letting text-to-image models reproduce a more accurate distribution of corresponding user prompts without additional manual engineering.
Our work makes the following key contributions:
@@
    We introduce an universal prompt optimization framework that is designed to work with any text-to-image model, providing a versatile solution for improving image generation across different architectures.
@@
    We provide comprehensive evaluation methodology and results including using both standard and non-standard metrics. Our experiments demonstrate that TIPO’s prompt engineering method can improve image quality in terms of mathematical measures and user preferences, validating its effectiveness from multiple perspectives.
Figure 1: Comparison of prompt optimization methods. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
This paper presents TIPO’s theoretical foundations (Section 3), implementation details (Section 4), and experimental results (Section 5) demonstrating its effectiveness in improving aesthetic scores [19], reducing image corruption [43, 44], and better aligning generated images with dataset distributions [28, 46]. Our findings highlight the critical role of prompt optimization in advancing T2I technology and open avenues for broader applications of automatic prompt refinement.
2 Related Works
2.1 Direct Prompt Refinement Through LLMs
With the advent of advanced LLMs, several projects have leveraged their few-shot in-context learning capabilities to optimize prompts for Text-to-Image (T2I) models. For instance, CogView3 [67] and the work by Lee et al. [32] utilize GPT-J and Text Style Transfer (TST) techniques for prompt refinement. However, this approach relies heavily on the LLM’s inherent knowledge of visual content descriptions, which can be unpredictable and may not always align with specific T2I model requirements.
2.2 Extending Prompts with Manually Collected Data
In the open-source T2I community, a common approach involves using curated datasets of “good prompts” to fine-tune or train LLMs for prompt extension [5, 60, 17]. These models are typically trained on image-text pairs collected from users of proprietary T2I services like Midjourney [^1] https://www.midjourney.com or Dall-E 3 [7]. While this method is straightforward and somewhat effective, it often lacks the versatility to adapt to diverse T2I models beyond the one used in the original dataset collection.
2.3 Image Output-Based Tuning
Recent works have explored more sophisticated prompt optimization techniques based on image outputs.
Reinforcement Learning Approaches: Promptist [26] introduced a two-stage framework combining supervised fine-tuning with reinforcement learning (RL). The RL stage incorporates aesthetic and relevance scores as environmental feedback, enabling the model to learn optimal prompt refinement strategies.
Iterative Optimization: OPT2I [41] proposed a backpropagation-free prompt optimization framework. This method uses scorers and LLMs to iteratively update input prompts, aiming to improve aesthetic quality and coherence.
These RL-based or RL-like approaches allow for prompt refinement tailored to specific model properties or user preferences. However, they often require extensive computational resources and may not generalize well across different T2I models or datasets.
While these methods have shown promise, they often face challenges such as difficulty in maintaining alignment with original user intent, challenges in collecting manually engineered prompts, high computational costs for RL-based methods, and limited generalization across diverse T2I models. By contrast, TIPO offers an efficient approach to prompt optimization by directly leveraging the training dataset, removing the need for manual prompt engineering and runtime tuning, and achieving versatile, contextually aligned image generation.
@@ 2.3 Image Output-Based Tuning
3 Prompt Optimization
3.1 Motivation and Challenges
Nowadays text-conditioned models [8, 48, 21] generate outputs based on user-provided text inputs. However, when users seek high-quality outputs or wish to explore diverse outputs while preserving key aspects of their input, they often need to manually expand their prompts. This manual prompt expansion can be time-consuming and labor-intensive.
To address these challenges, previous works have aimed to automate the prompt optimization process. Some methods attempt to expand prompts using the familiar vocabulary and stylistic tendencies of the model [25], while others rely on direct feedback from the model to iteratively improve prompt alignment [26]. However, these approaches have limitations: if the structure or language does not fully align with the model’s prompting conventions, performance may degrade, and feedback-based approaches often involve significant computational costs.
Therefore, effective prompt optimization for text-conditioned models requires accurately capturing user intent and expanding the prompt in a way that is both precise and aligned with the model’s conventions for interpreting text input. This need for precision and alignment motivates a systematic approach to prompt formulation, which we define in the following section.
3.2 Formulation
To formalize prompt extension, we define $\mathcal{P}$ as the set of all possible prompts, $\mathcal{N}$ as the Gaussian noise vector space, and $\mathcal{I}$ as the set of all possible images. A T2I model then acts as a mapping function:
$f(p):\mathcal{N}\rightarrow\mathcal{I}_{p}$
where, for any prompt $p\in\mathcal{P}$, the model maps noise vectors from $\mathcal{N}$ to a subset of images $\mathcal{I}_{p}\subseteq\mathcal{I}$ associated with the prompt $p$.
Now we define a pair of prompts: simple prompt $p_{s}$ and detailed prompt $p_{d}$. Formally, for a simple prompt $p_{s}$ and a detailed prompt $p_{d}$ that extends $p_{s}$, we have the following relationship:
$f(p_{d})(\mathcal{N})\subseteq f(p_{s})(\mathcal{N})$
This formulation implies that the image subset generated by the detailed prompt $p_{d}$ is contained within the subset generated by the simple prompt $p_{s}$, thereby preserving the relevance of the initial prompt while adding specificity.
To further encapsulate the prompt generation process, we define the prompt extension function, $\mathcal{E}$. This function takes a simple prompt $p_{s}$ and additional metadata $\mathcal{M}$ (such as tags) as inputs, and produces a detailed prompt $p_{d}$ as output:
$\mathcal{E}:\mathcal{P}_{s}\times\mathcal{M}\rightarrow\mathcal{P}_{d}$
where $\mathcal{P}_{s}$ is the space of simple prompts and $\mathcal{P}_{d}$ is the space of detailed prompts. Here, $\mathcal{M}$ provides auxiliary information that guides the extension of $p_{s}$ into a more refined prompt $p_{d}$.
For each instance of prompt extension, we apply the function as follows:
$p_{d}=\mathcal{E}(p_{s},M)$
where $M$ could represent a specific metadata component, such as $T_{s}$, or other forms of contextual information depending on the input type.
The simplest approach to extend a simple prompt is by adding random words. However, this approach has limitations, as it often introduces internally conflicting information that results in out-of-distribution (OOD) prompt inputs, leading the model to generate nonsensical images or the model to generate not user intent images. In such cases, although each added word may be meaningful in isolation, these random combinations can create prompts that are internally contradictory, as shown in Figure 2 (c). For example, the generated images may deviate from the user’s intent by not depicting an astronaut (left top), producing a horse-shaped tank (right top) instead of a horse, omitting Mars (left bottom), or even including a random snake (right bottom), which introduces nonsensical elements. To represent this random extension method, let $p_{r_{i}}$ denote the $i^{th}$ prompt created by adding random words to $p_{s}$. The resulting union of distributions fails to approximate the desired distribution.
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\supseteq\mathcal{I}\supseteq\mathcal{I}_{p_{s}}$
Which means:
$\bigcup_{i=0}^{n}f(p_{r_{i}})(\mathcal{N})\not\approx\mathcal{I}_{p_{s}}$
This scenario is illustrated in Figure 2 (c), where “Simple Prompt + Random Words” shows a pointy distribution extending beyond the bounds of the desired distribution or even the “image” distribution itself.
4 TIPO Framework and Model
4.1 Distribution-Aligned Prompt Extension
Central to TIPO’s method is an extension function $\mathcal{E}$, which iteratively refines prompts to enhance both the diversity and quality of generated images while remaining aligned with the intent of the original user-provided prompt and the training data distribution. By applying $\mathcal{E}$ to a simple prompt $p_{s}$ to generate a detailed prompt $p_{d}$, TIPO aims to approximate the target distribution:
$\bigcup_{i=0}^{n}f(p_{t_{i}})(\mathcal{N})\approx\mathcal{I}_{p_{s}}$
where $p_{t_{i}}$ denotes the $i^{th}$ TIPO-generated prompt derived from $p_{s}$, and $\mathcal{I}_{p_{s}}$ represents the target distribution of image outputs associated with $p_{s}$. Through this systematic prompt refinement, Figure 2 (d) “Simple Prompt + TIPO” illustrates the broader image space exploration enabled by TIPO while retaining relevance to the original prompt, effectively balancing diversity with coherence.
4.2 Prompt Structure
The construction of simple prompts, $p_{s}$, and their corresponding detailed prompts, $p_{d}$, varies depending on the dataset and input modality. We consider two primary scenarios: tag-based captions and natural language (NL) captions.
Tags are used to describe the content of images and are commonly formatted as a comma-separated string when input into T2I models. For a given image-tag pair, let $T_{n}=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ represent the complete set of tags for an image. We derive the simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ where $m<n$, and define the detailed tag set as $T_{d}=T_{n}$. The simple and detailed tag prompts are then constructed as:
$p_{s}=\text{concat}(T_{s}),\quad p_{d}=\text{concat}(T_{d})$
NL captions are typically generated by LMs with vision modality [36, 1, 2, 33, 6, 16, 63, 18] and may contain multiple sentences. Each image is associated with either two types of NL captions—a short NL caption $S_{s}$ providing a brief, general description, and a long NL caption $S_{d}$ offering an extended, detailed description—or a single long caption that can be segmented. For images with a single long caption, let
$S_{d}=\{\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}\}$
represent the complete set of sentences in $p_{d}$. We derive the simple sentence set
$S_{s}=\{\text{sentence}_{1},\text{sentence}_{2},\ldots,\text{sentence}_{m}\}$
where $m<n$, and define the detailed sentence set as $S_{d}=S_{n}$. The simple and detailed NL prompts are then constructed as:
$p_{s}=\text{concat}(S_{s}),\quad p_{d}=\text{concat}(S_{d})$
These ensure that $p_{d}$ is a superset of $p_{s}$, aligning with the concept that detailed prompts provide a more specific description.
To generate $p_{d}$ from $p_{s}$, TIPO employs two distinct formats based on the relationship between $p_{s}$ and $p_{d}$:
  1. 1.
    When $p_{s}$ is a substring of $p_{d}$: The format <meta> <p_d> is used. As TIPO model is an autoregressive LM, $p_{d}$ is directly used after the <meta> tokens, allowing the model to generate subsequent tokens based on any substring of $p_{d}$.
  2. 2.
    When $p_{s}$ is not a substring of $p_{d}$: The format <meta> <p_s> <p_d> is employed. Since $p_{s}$ is not contained within $p_{d}$, both are included separately after the <meta> tokens to guide the generation process effectively.
4.3 Prompt Formatting
In TIPO, we define a simple and consistent format for the content of the <meta> token or the simple prompt $p_{s}$. The format is as follows: <Category>: <Content>. Metadata categories might include artist, copyright, aspect ratio, quality, and year. For instance,
quality: masterpiece, year: newest
Importantly, in certain cases, input tag or NL prompts can be treated as metadata <Content> as well, when the task involves generating tags or NL prompts conditioned on each other.
4.4 Task Definitions

### v3

@@ Abstract
TIPO (Text-to-Image Prompt Optimization) introduces an efficient approach for automatic prompt refinement in text-to-image (T2I) generation. Starting from simple user prompts, TIPO leverages a lightweight pre-trained model to expand these prompts into richer, detailed versions. Conceptually, TIPO samples refined prompts from a targeted sub-distribution within the broader semantic space, preserving the original intent while significantly improving visual quality, coherence, and detail. Unlike resource-intensive methods based on large language models (LLMs) or reinforcement learning (RL), TIPO provides computational efficiency and scalability, opening new possibilities for effective, automated prompt engineering in T2I tasks.
We provide visual results, human preference report to investigate TIPO’s effectiveness. Experimental evaluations on benchmark datasets demonstrate substantial improvements in aesthetic quality, significant reduction of visual artifacts, and enhanced alignment with target distributions along with significant human preference proficiency. These results highlight the importance of targeted prompt engineering in text-to-image tasks and indicate broader opportunities for automated prompt refinement.
@@ 1 Introduction
The rapid proliferation of Text-to-Image (T2I) generative models have revolutionized artistic creation [48, 7, 20, 57, 53, 54, 60, 55, 50, 59, 15, 14, 35, 21, 8]. These models offer direct control over generative visual content via text prompts. To achieve precise control, modern T2I architectures are often trained on lengthy, detailed text descriptions, which may consist of individual, formatted tags of objects, backgrounds, styles, or complex, integrated paragraphs outlining image content and layout. However, the increasing complexity of prompts also forces users to iteratively refine their inputs, making high-quality T2I artwork accessible only to those with significant prompt engineering expertise.
Recently, efforts have been made to reduce the reliance on human expertise through prompt optimization, i.e., expanding and refining a user’s primitive input into a more detailed prompt to enhance generation quality. A straightforward approach is to leverage pre-trained Large Language Models (LLMs) to rewrite prompts in a zero-shot manner [42]. Yet, LLMs are primarily trained on general natural language, such as paragraphs and dialogues, which differ significantly from the structured prompts used for T2I models. This discrepancy often leads to additional effort in crafting LLM prompts and increased misalignment between generated images and intended prompts. A more effective approach is to train LLMs directly on prompt data collected from model users [5, 17]. While promising, this method is inherently constrained by the varying levels of user expertise, often resulting in inconsistent or suboptimal outputs. Recent work [26] trains LLM with reinforcement learning, where aesthetic scores of generated images serve as rewards. However, the reinforcement learning is performed on one specific T2I model with high computational cost, hindering its application to a broader variety of models.
Figure 1: Comparison of prompt optimization methods using LLM. (a) uses instructions for prompting but its understanding is constrained by the LLM’s knowledge base, not T2I model. (b) relies on a curated prompt database, enhancing detail but limiting variety by not fully leveraging the T2I model’s learned distribution. (c) optimizes using the scorer with RL, requiring multi-turn inference with additional cost. (d) aligns prompts with the T2I model’s training distribution, ensuring detailed and various topic-related prompt generation that fits the target T2I model.
Deviating from existing work, we argue that the optimal prompt lies in T2I models’ training set. Precisely, when a prompt aligns with the data distribution used for model training, the model can better interpret the prompt, resulting in improved text alignment and higher image quality. Based on this insight, we introduce TIPO (Text to Image with text pre-sampling for Prompt Optimization), an innovative framework designed to enhance T2I generative models. At its core lies a lightweight language model trained on multiple meticulously designed pretext tasks that transform prompts into tags or sentences from coarse to fine. Leveraging the trained model, TIPO can progressively and seamlessly adapt to various user input types, generating accurate and robust prompts for various T2I models. In addition, our training data is derived from a curated text description dataset comprising over 30 million text-image pair samples and more than 20 billion tokens, ensuring comprehensive coverage of T2I training sets. Extensive experiments on test prompts from in- and out-of-domain demonstrate that TIPO consistently outperforms state-of-the-art methods in fidelity, aesthetics, and text alignment. Detailed analysis shows that TIPO has up to 62.80% win rate in user preference and 25% improvement in runtime efficiency over the second-best baseline. To summarize, our contributions are at least three-fold:
@@
    We introduce TIPO, a prompt optimization framework that leverages the large-scale text data distributions used in text-to-image (T2I) training.
@@
    We train a multi-task language model, which adaptively and progressively refines user inputs from either tag-based or natural language prompts into unified formats, enhancing compatibility across a broad spectrum of T2I models.
  3. 3.
    Extensive experiments demonstrate that TIPO achieves superior image quality, strong text alignment, higher human preference, and improved runtime efficiency across both in-domain and out-of-domain prompts, highlighting its great potential in real-world applications.
2 Related Work
Prompt optimization for T2I models typically leverages language models to refine user inputs. They can be broadly classified into two categories: (1) Model-specific strategies that tailor prompts for a particular T2I model, and (2) Universal strategies that aim to improve prompt quality across a variety of T2I models.
2.1 Model-specific Strategies
T2I models generate images whose quality is often measured using metrics such as fidelity, aesthetics, and user preference. These metrics facilitate reinforcement learning approaches that optimize prompts for a specific T2I model. For instance, Promptist [26] fine-tunes a pre-trained language model by using CLIP relevance scores as rewards. Similarly, PAE [44] extends this approach by generating dense text embeddings rather than discrete text tokens, with additional control vectors during online reinforcement learning. However, these methods are computationally intensive, often struggling with a larger number of training prompts. Moreover, a model optimized for one specific T2I system may not generalize well to others. In contrast, our method leverages over 30 million text descriptions to cover a wide range of high-quality prompts, ensuring compatibility with a broad spectrum of T2I models.
2.2 Universal Strategies
To reduce the dependency on specific T2I models, some researchers have focused on refining prompts solely using language models. For example, CogView3 [71] employ GLM-4 [24], and Lee et al. [33] employ GPT-J and Text Style Transfer (TST) techniques, respectively, for prompt enhancement. However, both of them rely heavily on the LLM’s inherent understanding of visual content descriptions, which may result in a misalignment with the diverse requirements of various T2I models. Alternatively, other approaches collect high-quality prompts from T2I model users to fine-tune or train LLMs [5, 63, 17]. Such methods, however, are limited by the inconsistent expertise of users. Conversely, our approach constructs both tag-based and natural language prompts using a large-scale dataset of image-text descriptions, thereby aligning closely with the text distributions underlying T2I models.
3 Preliminaries
Notations. We present the notations in our work:
  • $\mathcal{P}$: The space of all possible prompts.
  • $\mathcal{I}$: The space of all possible images.
  • $p$: User-provided prompt.
  • $p_{o}$: Optimized prompt.
  • $p_{ps}$: Sampled modified prompt derived from p.
  • $\mathcal{M}$: Metadata (e.g., style tags) aiding prompt transformation.
  • $\mathcal{T}$: Transformation function for pre-sampling.(e.g., Random augmentation, TIPO or LLM)
  • $T_{p},T_{ps}$: Sets of tags for original and pre-sampled prompts.
  • $S_{p},S_{ps}$: Ordered lists of natural language descriptions for original and pre-sampled prompts.
Text-to-Image Model. A text-to-image (T2I) model is a mapping function
$f(p)\;\;\to\;\;\mathcal{I}_{p}\subseteq\mathcal{I},$
which maps a user-provided text prompt $p$ to a subset of images $\mathcal{I}_{p}$ corresponding to the prompt.
Problem statement. Let $\mathcal{I}_{u}$ denote the user’s intended distribution over images. The task of Prompt optimization is to generate a prompt $p_{o}$ that minimizes a distance $d$ between the T2I output distribution $\mathcal{I}_{p}$ and $\mathcal{I}_{u}$:
$p_{o}\;=\;\arg\min_{p\in\mathcal{P}}\;d\bigl{(}\mathcal{I}_{p},\,\mathcal{I}_{u}\bigr{)},$
where $d$ is a suitable metric (e.g., Frechet Inception Distance) comparing two image distributions.
@@ 3 Preliminaries
4 Methodology
We aim to optimize user prompts to enhance image generation quality. Instead of end-to-end optimizations tailored to a single T2I model, our focus is on prompt rewriting that generalizes across a broad spectrum of models. Our core intuition is that an ideal prompt should align with the texts used in T2I model training. However, rapid advances in image captioning have rendered these texts increasingly diverse and complex. To address this, we propose to (1) design a clearly structured prompt schema compatible with most text descriptions, and (2) implement a pre-sampling algorithm that progressively refines arbitrary, coarse user input into organized, fine-grained prompts.
4.1 Text Set Preparation
Although the text descriptions used for T2I model training are notably diverse, most are image captions that fall into two broad categories: tag-based and natural language (NL)-based captions. Tag-based captions, such as those in the Danbooru2023 dataset [46, 68], use comma-separated, succinct terms to describe image content. In contrast, NL-based captions, typically generated by language models with visual capabilities [37, 1, 2, 34, 6, 16, 67, 18], may comprise multiple sentences. We represent both types using a unified text set $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$, where each element is an individual tag or sentence.
While the original image captions are fine-grained and detailed, which can yield high-quality images when all elements are used, they often result in prompts that are excessively lengthy or overloaded with information. Such prompts diverge from typical user input and pose alignment challenges. To mitigate this, we construct a simpler subset $T_{s}\subset T$ by removing some tags and sentences from the original set [^1] For long sentences, off-the-shelf text splitters (e.g., FastText) are used to segment them into multiple sentences., as detailed in Section 4.2.
4.2 Formatted Prompt Construction
We aim to construct prompts in a unified format compatible with existing image captions. First, we incorporate the common metadata present in these image captions, typically represented as <Category>: <Content>. These metadata categories primarily include artist, copyright, aspect ratio, quality, and year (e.g., quality: masterpiece, artist: Picasso). This structured metadata is intuitive for users to read and edit, while also providing strong guidance to downstream T2I models on the generation scope.
Next, we construct both tag-based and NL-based prompts using text sets $T$. Our design generates both simple (incomplete) and complete prompts for each image, and we train an auto-regressive language model to extend the simple prompts into complete versions. For tag-based prompts, since the tags are largely order-insensitive (i.e., the order has minimal impact on T2I outcomes), we propose a prefix-based dropout strategy. We first randomly shuffle the complete set of tags $T=\{t_{1},t_{2},t_{3},\ldots,t_{n}\}$ from a given image caption. Then, we construct a simple tag set $T_{s}=\{t_{1},t_{2},\ldots,t_{m}\}$ by randomly selecting $m<n$ tags. The prompts are constructed as:
$p_{s}=\text{concat}(T_{s}),\quad p_{o}=\text{concat}(T)$
Because $p_{s}$ is always a prefix of $p_{o}$, the language model can readily expand the simple tag-based prompt into its complete version.
For NL-based prompts, however, this strategy cannot be applied directly because the first sentence often contains crucial information [25] and the order of sentences significantly influences the caption’s semantics. Therefore, we preserve the first sentence and randomly drop some of the subsequent sentences without changing their order. Let:
$S=[\text{sentence}_{1},\text{sentence}_{2},\text{sentence}_{3},\ldots,\text{sentence}_{n}]$
represent the ordered sequence of sentences in an image caption. We derive a simple subsequence $S_{s}$ by randomly selecting $m$ sentences from $S$ while ensuring that the first sentence is always included and that the original order is maintained. In other words,
$S_{s}=[\text{sentence}_{1},\text{sentence}_{i_{2}},\ldots,\text{sentence}_{i_{m}}],$
with $1<i_{2}<\ldots<i_{m}\leq n$ and $m<n$. The simple and complete NL-based prompts are then constructed as:
$p_{s}=\text{concat}(S_{s}),\quad p_{o}=\text{concat}(S_{s},S)$
This ensures that $p_{s}$ remains a prefix of $p_{o}$. Although some sentences may be repeated in $p_{o}$, selecting a smaller $m$ effectively mitigates this, and it does not empirically affect the generation quality.
4.3 Text Pre-sampling
We aim to align user input with the constructed high-quality training prompts $p_{o}$ via pre-sampling, which stands for “text sampling before image sampling” - a process that first constrains the target distribution in text space before proceeding to image generation. Conceptually, while pre-sampling strategies generally add more text tokens to user input, the outcomes of different strategies may still vary (as illustrated in Figure 2).
To enrich the image content while mitigating undesired output, we perform pre-sampling based on a language model trained on $p_{o}$. A straightforward pre-sampling strategy is to do text completion, i.e., directly append output tokens to the end of user input. However, we posit that the user input is unlikely to be a part of a good prompt. The na’́ive text-completion strategy prioritizes the user’s original input, resulting in worse outcomes as evidenced by Section 5.6.